Distance =14,23,9,17,10,22,5,12,6,10. Charges= 68,105,40,79,81,95,31,72,45,93.
Required following.
1) coefficient of correlation
2) coefficient of determination
3) probable Error
4) Interpretation
1
Expert's answer
2020-05-07T19:19:56-0400

1) X Values: ∑ = 128,Mean = 12.8

∑(X - Mx)2 = SSx = 345.6

Y Values: ∑ = 709,Mean = 70.9

∑(Y - My)2 = SSy = 5626.9

X and Y Combined

N = 10

∑(X - Mx)(Y - My) = 1113.8

r = ∑((X - My)(Y - Mx)) / "\sqrt{((SSx)(SSy))}"

r = 1113.8 / "\sqrt{((345.6)(5626.9))}" = 0.7987

2) cd=r2=0.63793

3)using formula (where 0.674 - const): e=0.674 *(1-r2) /"\sqrt{N}" =0.07717

​4)A strong uphill (positive) linear relationship